OpenStudy (anonymous):

Write the equation of a line in slope-intercept form that has a slope of -2 and passes through the point (2, -8).

OpenStudy (anonymous):

y = A x + B A = slope = -2 y = -2 x + B now sub the point -8 = -2 * 2 + B -8 = -4 + B B = -4 y = -2x -4

OpenStudy (dumbcow):

 thats the idea, now you can also plug it into a formula \[y-y_1 = m(x-x_1)\] \[y+8 = -2(x-2)\] \[y =-2x -4\]